Joss is x years older than Noah . If Noah is 13 years old , how old is joss
Answer by jim_thompson5910(35256)


Answer: x+13
This is the same as 13+x

Explanation:
Noah is 13 years old. Add on x more years and you get 13+x or x+13, which is Joss's age. If we knew the value of x, then we could evaluate the expression. For instance, if x = 2, then 13+x = 13+2 = 15. However, we don't know what x is, so we leave the expression as is.